# Towards Mitigating Excessive Forgetting in LLM Unlearning via Entanglement-Aware Unlearning with Proxy Constraint

### v1

@@ 5.2 EAGLE with Proxy Constraint (EAGLE-PC)
To this end, we propose EAGLE-PC, an extension of EAGLE that incorporates Proxy Constraint based on test inputs synthesized through in-context learning (ICL) with LLM. These proxy samples emulate how the model would naturally respond to forget samples if it has never been exposed to them. By applying a loss-based soft constraint on the model’s predictions over these proxy inputs during unlearning, EAGLE-PC provides a semantically grounded and adaptive boundary that not only eliminates the need for manual threshold tuning, but also effectively mitigates over-forgetting and preserves the model’s ability to generalize after unlearning.
@@
Unlike traditional test data, which typically relies on real, unseen ground-truth labels, our approach introduces novel unlearning constraints by leveraging a language model to generate plausible reference answers, referred to as test answers. These test answers are generated via In-Context Learning (ICL), where the language model is prompted with carefully selected exemplars that contrast correct outputs with uninformed responses. This design enables the ICL-generated outputs to serve as high-quality surrogates that approximate the behavior of the retrain model on unseen inputs. As semantically aligned references, they offer an effective criterion for identifying over-forgetting in forget samples. An additional advantage of employing ICL-based proxy answers lies in their adaptability to specific model behaviors and data distributions. By carefully selecting in-context exemplars that mirror the target model’s response tendencies, both in terms of correct predictions and typical failure cases, the language model can simulate how the target model would behave when presented with similar inputs after unlearning. This enables the generated test data to more accurately reflect model-specific generalization under ignorance, resulting in a penalty signal better aligned.
@@
During unlearning, we monitor the prediction loss of each forget sample with respect to its original label and compare it to the loss computed using the LLM-generated proxy answer. If the loss on the real answer surpasses that on the proxy baseline, we impose a penalty term that discourages excessive divergence from the model’s natural generalization capability. This mechanism constrains the unlearning process to remain aligned with the behavior of the retrain model, serving as an early warning signal to prevent over-forgetting.
@@
Algorithm 3 Generating Proxy Answers and Computing Over-Forgetting Penalty $\mathcal{P}(\xi_{t},\mu)$
@@
9:Step 2: Compute Over-Forgetting Penalty
@@
The detailed procedure is shown in Algorithm 3. In each iteration, given forget samples $\xi_{t}$, each sample $z_{i}=\{x_{i},y_{i}\}$ consists of a question $x_{i}$ and its ground-truth answer $y_{i}$. We first prepare few-shot exemplars $\{(x_{j},y_{j},y_{j}^{\prime})\}$ to prompt a pretrained large language model (LLM). Here, $y_{j}^{\prime}$ denotes the model’s expected answer to $x_{j}$ without ever having seen the forget data. Conditioned on these exemplars, the LLM synthesizes a proxy answer $y_{i}^{test}$ for the target question $x_{i}$. The resulting pair $z_{i}^{test}=\{x_{i},y_{i}^{test}\}$ serves as a test reference approximating the model’s natural prediction behavior under the absence of explicit memorization. All proxy samples are collected to form the test dataset $\mathcal{D}_{test}$. Next, for each forget sample $x_{i}$, we compute its current loss $f(\theta_{t},z_{i})$ and the loss of proxy data $f(\theta_{t},z_{i}^{test})$. If the model’s loss on the forget data drops below the loss of proxy data $f(\theta_{t},z_{i})<f(\theta_{t},z_{i}^{test})$, this indicates that the unlearning process may have overshot the target, deviating excessively from the model’s expected generalization. To softly penalize such over-forgetting, we accumulate a penalty $\mathcal{P}$ defined as: $\mathcal{P}\leftarrow\mathcal{P}+\mu\cdot(f(\theta_{t},z_{i}^{test})-f(\theta_{t},z_{i}))\quad\text{if}\quad f(\theta_{t},z_{i})<f(\theta_{t},z_{i}^{test})$, where $\mu$ is a hyper-parameter controlling the penalty’s strength. This penalty is then added to the overall unlearning objective.
@@
In Algorithm 3, the examples used in ICL are formatted as question–answer (QA) pairs, following the convention established by the TOFU benchmark. This approach can also be extended to datasets that are not originally in QA format by converting their content into QA pairs and reverting them back to the original form. While Algorithm 3 presents the process as querying the LLM once per sample, in practice we adopt a more efficient strategy by batching multiple forget samples into a single prompt. This allows the LLM to generate corresponding proxy answers in parallel, substantially reducing the computational overhead and enhancing scalability to larger unlearning datasets. Our method is not restricted to QA datasets and can be naturally extended to other data modalities. For instance, narrative passages can be reformulated into QA pairs by framing key facts as questions and answers, while dialogue data can be adapted by converting turns into query–response pairs.
We present a representative example below, which includes a user query, the corresponding real (gold-standard) answer, the answer given by the model to the question without learning the corresponding knowledge, and the test answer generated via in-context learning (ICL) using GPT-4o. Notably, both the retrain and test answers exhibit similar response patterns and share comparable deviations from the real answer. This alignment suggests that the test answers can effectively approximate the retrain model’s behavior, capturing how forgetting may influence the outputs.
@@
Table 1: PHI Results on TOFU benchmarks with different forgetting ratios (1%, 5%, and 10%). Each method is evaluated using four metrics: F.Q., M.U., and R-L represent forget quality, model utility and ROUGE-L respectively. For each baseline (GA, NPO, GD, NPO+GD), we report its original performance and the corresponding results when enhanced with EAGLE-PC.
@@
F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$ F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$ F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$
@@
EAGLE-PC(GA) 0.9999 0.3260 0.4223 0.4843 1.12e-5 0.4094 0.4261 0.4064 2.36e-4 0.3399 0.4106 0.3505
EAGLE-PC(NPO) 0.9900 0.3016 0.4133 0.4762 1.12e-5 0.3962 0.4323 0.4042 3.77e-5 0.3718 0.4366 0.3892
EAGLE-PC(GD) 0.9900 0.3594 0.4502 0.5564 1.83e-5 0.3520 0.4327 0.4181 5.52e-5 0.3356 0.4363 0.3925
EAGLE-PC(NPO+GD) 0.9900 0.3366 0.4610 0.5558 0.7934 0.3474 0.4259 0.4147 0.3958 0.3601 0.4209 0.4211
6 Experiments
In this section, we comprehensively evaluate our proposed unlearning framework EAGLE-PC by applying it to a widely adopted benchmark setting for LLM unlearning: the removal of knowledge related to a fictional author from the TOFU [18] and the mitigation of memorized or privacy-sensitive content in the MUSE [22] benchmark. Furthermore, in Section 6.3.1, we present a detailed comparison between our entanglement-aware loss reweighting strategy and prior unlearning approaches, including uniform baselines, empirical-loss-based heuristics [10] and the saturation-importance reweighting strategy SatImp [28], to highlight the advantages of structurally informed coefficient estimation. Details of the experimental setup and metrics are provided in Appendix A.2.
Figure 3: Loss difference on Forget 1% between unlearned models and retrain model for each forget/retain sample using the Phi model. EAGLE-PC yield smaller loss gaps, indicating closer alignment with retraining and improved forgetting precision with minimal utility degradation.
Table 2: Performance of unlearning methods on TOFU benchmark using LLaMA2-7B across different unlearning ratios (1%, 5%, 10%). We report Forget Quality (F.Q.), Retain Language modeling score (R-L), and Model Utility (M.U.) on both forget and retain sets. EAGLE-PC variants achieve consistently better performance than baselines across all settings.
Method TOFU-1% TOFU-5% TOFU-10%
F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$ F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$ F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$
@@ 6 Experiments
EAGLE-PC(GA) 0.9188 0.3719 0.5660 0.7762 1.12e-5 0.4808 0.5106 0.5112 1.16e-5 0.0018 0 0.0040
EAGLE-PC(NPO) 0.9188 0.3682 0.5654 0.7733 1.12e-5 0.4504 0.5119 0.4912 2.17e-7 0.4987 0.5178 0.5176
EAGLE-PC(GD) 0.7659 0.3391 0.5325 0.6814 2.38e-6 0.0621 0.6213 0.5645 0.0043 0.0761 0.6189 0.6659
EAGLE-PC(NPO+GD) 0.9188 0.2417 0.6454 0.8059 0.9647 0.3022 0.6202 0.7926 0.3417 0.2735 0.6315 0.7266
6.1 Experiments on TOFU
@@
We evaluate our unlearning method on the TOFU benchmark [18], which tests the removal of factual knowledge about 200 fictitious authors, each with 20 QA pairs. TOFU provides three forget sets $\mathcal{D}_{f}$ (1%, 5%, 10% of authors), while the retain set $\mathcal{D}_{r}$ contains QA pairs of the remaining authors. Forget quality [18] measures how closely the unlearned model matches one trained only on $\mathcal{D}_{r}$, whereas model utility reflects performance on held-out retain data regarding fictional writers, real-world writer profiles, and other world facts. We also report ROUGE-L for both forget and retain performance. Experiments use fine-tuned Phi-1.5 [15] and LLaMA2-chat-7B [25], both pretrained on all 200 authors. We report experimental results on the TOFU benchmark using the Phi model in Table 1. The table includes comparisons across a range of unlearning methods: Gradient Ascent (GA), Gradient Difference (GD), Negative Preference Optimization (NPO), and NPO+GD which denotes NPO on the forget set while simultaneously conducting gradient descent on the retain set, each both in vanilla form and integrated with EAGLE-PC.
@@
Consistent Performance Gains Across Metrics. Our framework consistently outperforms the vanilla baselines across both Forget Quality and Model Utility. For instance, under the 5% forgetting setting, integrating NPO+GD into our framework significantly elevates the F.Q. from 0.0878 to 0.7934, while simultaneously enhancing the model utility from 0.4158 to 0.4259. Moreover, the Rouge-L score on the retain set improves from 0.3942 to 0.4147, demonstrating a superior balance between unlearning accuracy and model performance.
@@
Consistent Gains Across Optimizers. Such consistent improvements are observed across all optimizers considered (e.g., GA, GD, NPO, and NPO+GD). Although GA performs less favorably than NPO, and GD underperforms compared to NPO+GD, the integration of all optimizers into our framework leads to noticeable gains. Notably, even though the trade-off between F.Q. and M.U. for GA under our framework remains inferior to the vanilla NPO baseline, it still benefits from our enhancements. We attribute this to the limited potential of the GA optimizer itself. In contrast, the combination of NPO+GD within our framework (denoted as EAGLE-PC(NPO+GD)) achieves a forgetting quality that closely approximates that of the retrain model, demonstrating our framework’s strong capability in precise and efficient unlearning.
Retain Dataset Enhances Unlearning Performance. It can be observed that unlearning methods which incorporate the retain dataset during optimization achieve clearly better trade-offs between forgetting and utility. Specifically, GD outperforms GA across all forgetting ratios—for example, under 1% forgetting, GD achieves higher model utility (0.4564 vs. 0.4151) and ROUGE-L (0.5456 vs. 0.4722) while maintaining strong forget quality. Similarly, NPO+GD consistently surpasses NPO, with improved utility (e.g., 0.4158 vs. 0.4203 at 5%) and robustness across metrics. These results highlight the value of incorporating retain data into the forgetting process to avoid over-forgetting and preserve generalization.
@@
EAGLE-PC framework requires only the average retain embedding yet outperforms methods that use the full retain dataset. The EAGLE-PC framework enables even weaker optimizers to outperform stronger baselines by incorporating lightweight, entanglement-aware guidance. A notable example is EAGLE-PC(GA), which builds upon GA and requires only the average embedding of the retain dataset. In contrast, GD depends on access to the full original retain dataset. Remarkably, EAGLE-PC(GA) consistently outperforms GD across multiple forgetting settings. Under the 5% unlearning setting, EAGLE-PC(GA) achieves a higher model utility (0.4261 vs. 0.4047) than GD, while also slightly improving forget quality (1.12e-5 vs. 1.39e-6). This demonstrates that even coarse-grained retain signals, such as a single average embedding vector, can provide strong regularization to guide forgetting without explicit access to the full retain dataset or gradient-level control. The same trend holds across multiple forgetting ratios. For instance, under 10% forgetting, EAGLE-PC(GA) boosts utility from 0.3878 to 0.4106, overtaking GD’s 0.4032 again, using only a minimal representation of the retain distribution.
These improvements can be primarily attributed to effectively addressing two major forms of over-forgetting in our framework: (1) Sample-level over-forgetting, where specific forget samples are excessively forgotten, leading to unnecessary utility degradation. By introducing Entanglement-awareness guided loss reweighting strategy, our method prevents disproportionate forgetting signals on individual samples and thus achieves more balanced unlearning. (2) Global over-forgetting, where the overall unlearning process drifts too far from the original model distribution, resulting in substantial performance loss on retain tasks. To mitigate this global risk, we incorporate a carefully designed proxy constraint that softly regularizes the unlearning trajectory, ensuring that knowledge removal remains controlled and utility-preserving.
To further validate the generality and robustness of EAGLE-PC, we conduct additional experiments using the LLaMA2-7B model on the TOFU benchmark under 1%, 5%, and 10% unlearning settings. As shown in Table 2, EAGLE-PC consistently outperforms baseline methods (GA, NPO, GD, and NPO+GD) across varying unlearning intensities in terms of both forget quality (F.Q.) and retain-side utility (measured via Model Utility and Retain LM score). For instance, under the 5% unlearning setting, EAGLE-PC(NPO+GD) boosts the forget quality from 7.54e-5 to 0.9647, while also improving model utility from 0.5905 to 0.6202, and enhancing the Retain ROUGE-L from 0.5681 to 0.7926. Notably, similar improvements are observed across other unlearning ratios. Under the 10% unlearning setting, EAGLE-PC(NPO+GD) achieves a threefold increase in forget quality compared to its base (0.3417 vs. 5.02e-10), while preserving substantially more useful knowledge, as evidenced by improvements in retain ROUGE-L (0.7266 vs. 0.4815) and model utility (0.6315 vs. 0.5608). To disentangle the contributions of individual components, we additionally report the performance of each component-injected variant in Appendix A.4.
@@
We further illustrate the effectiveness of EAGLE-PC by visualizing, in Figure 3 the per-sample absolute loss gap between each unlearned model and the retrain model on both the forget and retain datasets. The top row shows results on the forget dataset, where we observe that models incorporating EAGLE consistently produce smaller loss gaps compared to their counterparts. EAGLE better approximate the behavior of retraining in removing memorized knowledge. The bottom row presents results on the retain dataset. Notably, methods built upon our entanglement-aware framework consistently yield loss patterns that more closely align with the retrain model, compared to their non-entanglement counterparts. This indicates that by assigning differentiated forgetting effort based on sample entanglement, our approach enhances forgetting precision while significantly mitigating undesired forgetting of the retain knowledge, thus achieving a more stable and utility-preserving unlearning process.

### v2

@@ 5.2. EGU with Proxy Constraint (EGUP)
To this end, we propose EGUP, an extension of EGU that incorporates Proxy Constraint based on test inputs synthesized through in-context learning (ICL) with LLM. These proxy samples emulate how the model would naturally respond to forget samples if it has never been exposed to them. By applying a loss-based soft constraint on the model’s predictions over these proxy inputs during unlearning, EGUP provides a semantically grounded and adaptive boundary that not only eliminates the need for manual threshold tuning, but also effectively mitigates over-unlearning and preserves the model’s ability to generalize after unlearning.
@@
Unlike traditional test data, which typically relies on real, unseen ground-truth labels, our approach introduces novel unlearning constraints by leveraging a language model to generate plausible reference answers, referred to as test answers. These test answers are generated via In-Context Learning (ICL), where the language model is prompted with carefully selected exemplars that contrast correct outputs with uninformed responses. This design enables the ICL-generated outputs to serve as high-quality surrogates that approximate the behavior of the retrained model on unseen inputs. As semantically aligned references, they offer an effective criterion for identifying over-unlearning in forget samples. An additional advantage of employing ICL-based proxy answers lies in their adaptability to specific model behaviors and data distributions. By carefully selecting in-context exemplars that mirror the target model’s response tendencies, both in terms of correct predictions and typical failure cases, the language model can simulate how the target model would behave when presented with similar inputs after unlearning. This enables the generated test data to more accurately reflect model-specific generalization under ignorance, resulting in a penalty signal better aligned.
@@
During unlearning, we monitor the prediction loss of each forget sample with respect to its original label and compare it to the loss computed using the LLM-generated proxy answer. If the loss on the real answer surpasses that on the proxy baseline, we impose a penalty term that suppresses unlearning and prevents over-unlearning. This mechanism constrains the unlearning process to remain aligned with the behavior of the retrained model, serving as an early warning signal to prevent over-unlearning.
Figure 6. Comparison of unlearning performance on Phi‑1.5 and LLaMA2‑7B across Vanilla, SatImp (Yang et al., 2025), empirical‑loss‑based heuristics (Huang et al., 2024), EGUP, and full retraining. EGUP shows performance closest to full retraining.
@@
Algorithm 2 Generating Proxy Answers and Computing over-unlearning Penalty $\mathcal{P}(\xi_{t},\mu)$
@@
9:Step 2: Compute over-unlearning Penalty
@@
The detailed procedure is shown in Algorithm 2. In each iteration, given forget samples $\xi_{t}$, each sample $z_{i}=\{x_{i},y_{i}\}$ consists of a question $x_{i}$ and its ground-truth answer $y_{i}$. We first prepare few-shot exemplars $\{(x_{j},y_{j},y_{j}^{\prime})\}$ to prompt a pretrained large language model (LLM). Here, $y_{j}^{\prime}$ denotes the model’s expected answer to $x_{j}$ without ever having seen the forget data. Conditioned on these exemplars, the LLM synthesizes a proxy answer $y_{i}^{test}$ for the target question $x_{i}$. The resulting pair $z_{i}^{test}=\{x_{i},y_{i}^{test}\}$ serves as a test reference approximating the model’s natural prediction behavior under the absence of explicit memorization. All proxy samples are collected to form the test dataset $\mathcal{D}_{test}$. Next, for each forget sample $x_{i}$, we compute its current loss $f(\theta_{t},z_{i})$ and the loss of proxy data $f(\theta_{t},z_{i}^{test})$. If the model’s loss on the forget data drops below the loss of proxy data $f(\theta_{t},z_{i})<f(\theta_{t},z_{i}^{test})$, this indicates that the unlearning process may have overshot the target, deviating excessively from the model’s expected generalization. To softly penalize such over-unlearning, we accumulate a penalty $\mathcal{P}$ defined as: $\mathcal{P}\leftarrow\mathcal{P}+\mu\cdot(f(\theta_{t},z_{i}^{test})-f(\theta_{t},z_{i}))\quad\text{if}\quad f(\theta_{t},z_{i})<f(\theta_{t},z_{i}^{test})$, where $\mu$ is a hyper-parameter controlling the penalty’s strength. This penalty is then added to the overall unlearning objective.
@@
In Algorithm 2, the examples used in ICL are formatted as question–answer (QA) pairs, following the convention established by the TOFU benchmark. Our method can also be naturally extended to other data modalities. For instance, narrative passages can be reformulated into QA pairs by framing key facts as questions and answers, while dialogue data can be adapted by converting turns into query–response pairs. While Algorithm 2 presents the process as querying the LLM once per sample, in practice we adopt a more efficient strategy by batching multiple forget samples into a single prompt. This allows the LLM to generate corresponding proxy answers in parallel, substantially reducing the computational overhead and enhancing scalability to larger unlearning datasets. Below, we illustrate the prompt design used to obtain proxy answers through in-context learning (ICL). The prompt is structured into three components: System Instruction, which defines the model’s behavior; Input Description, which specifies the task context and information provided; and Output Requirement, which clarifies the expected form of the model’s response.
Prompt Template System Instruction: You are required to infer the pure_answer, meaning the answer a model would output without learning the relevant knowledge. Input Description: Each sample contains a question, the original answer, and a pure_answer produced by model that had never been exposed to the corresponding information. Your task is to analyze patterns of pure_answer and predict it for unseen questions. Output Requirement: Produce the most probable pure_answer rather than the true real-world answer.
We present an example below regarding fictitious author Basil Mahfouz Al-Kuwaiti, which includes the user query, the corresponding real (ground‑truth) answer, the pure answer produced by the model before acquiring the relevant knowledge, and the test answer generated via in‑context learning (ICL) using GPT‑4o. While the real answer identifies his genre as “French Literature”, the pure answer suggests “Alternate History”, and the ICL-based test answer produces “Science Fiction”. As Alternate History is widely recognized as a subgenre of Science Fiction, the test answer exhibits strong semantic alignment with the pure answer. Both outputs deviate comparably from the real answer, suggesting that the ICL approach effectively approximates the behavioral characteristics of the retrained model, capturing its essential response tendencies.
@@
6. Experiments
In this section, we comprehensively evaluate EGUP under widely adopted benchmark settings for LLM unlearning. Specifically, we follow two representative scenarios: (i) removing knowledge related to a fictional author in the TOFU benchmark (Maini et al., 2024), and (ii) mitigating memorized or privacy-sensitive content in the MUSE benchmark (Shi et al., 2024) (Section 6.3). Detailed experimental setups and evaluation metrics are provided in the appendix.
Table 1. Results on TOFU benchmarks with different forgetting ratios (1%, 5%, and 10%). Each method is evaluated using four metrics: F.Q., M.U., and R-L represent forget quality, model utility and ROUGE-L respectively.
@@ 6. Experiments
F.Q. $\uparrow$ R-L M.U. $\uparrow$ R-L F.Q. $\uparrow$ R-L M.U. $\uparrow$ R-L F.Q. $\uparrow$ R-L M.U. $\uparrow$ R-L
Phi-1.5
EGUP(GA) 0.9999 0.3260 0.4282 0.4764 1.12e-5 0.4094 0.4261 0.4064 2.36e-4 0.3399 0.4106 0.3505
EGUP(NPO) 0.9900 0.2946 0.4266 0.4723 1.12e-5 0.3962 0.4327 0.4082 5.52e-5 0.3847 0.4417 0.3983
EGUP(GD) 0.9900 0.2884 0.4567 0.5186 1.12e-5 0.3621 0.4391 0.4351 5.52e-5 0.3356 0.4363 0.3925
EGUP(NPO+GD) 0.9900 0.3366 0.4610 0.5558 0.7934 0.3474 0.4259 0.4147 0.3958 0.3601 0.4209 0.4211
@@
LLaMA2-7B
EGUP(GA) 0.9188 0.3719 0.5660 0.7762 1.12e-5 0.4808 0.5106 0.5112 1.16e-5 0.0018 0 0.0040
EGUP(NPO) 0.9900 0.3605 0.5648 0.7543 1.12e-5 0.4504 0.5119 0.4912 2.17e-7 0.4987 0.5178 0.5176
EGUP(GD) 0.7659 0.3391 0.5325 0.6814 2.96e-5 0.0000 0.6244 0.5694 0.0043 0.0761 0.6189 0.6659
EGUP(NPO+GD) 0.9188 0.2417 0.6454 0.8059 0.9647 0.3022 0.6202 0.7926 0.3417 0.2735 0.6315 0.7266
@@
6.1. Experiments Results
@@
We evaluate our unlearning method on the TOFU benchmark (Maini et al., 2024), which tests the removal of factual knowledge about 200 fictitious authors, each with 20 QA pairs. TOFU provides three forget sets $\mathcal{D}_{f}$ (1%, 5%, 10% of authors), while the retain set $\mathcal{D}_{r}$ contains QA pairs of the remaining authors. Forget quality (Maini et al., 2024) measures how closely the unlearned model matches one trained only on $\mathcal{D}_{r}$, whereas model utility reflects performance on held-out retain data, including information about fictional writers, real-world author profiles, and other world facts. Additionally, we report ROUGE-L scores for both forget and retain performance.
Experiments are conducted on fine-tuned Phi-1.5 (Li et al., 2023) and LLaMA2-chat-7B (Touvron et al., 2023), both pretrained on all 200 authors. Table 1 presents the experimental results on the TOFU benchmark, including comparisons across several unlearning methods: Gradient Ascent (GA), Gradient Difference (GD), Negative Preference Optimization (NPO), and their combination (NPO+GD), each reported in both vanilla form and when integrated with EGUP. The table highlights how EGUP consistently improves the vanilla baselines across different forgetting ratios, demonstrating enhanced forget quality and better retention stability. A broader comparison between EGUP and other state-of-the-art unlearning methods is illustrated in Figure 6.
Figure 7. Loss difference on Forget 1% between unlearned models and retrained model for each forget/retain sample using the Phi model. EGUP yield smaller loss gaps, indicating closer alignment with the retrained model.
Figure 8. Forget and retain loss on Forget-05 are compared under different GD variants against retraining. EGU(GD) drops the proxy constraint and EGUP_w/o(GD) drops the entanglement guidance. The results show that EGUP(GD) best balances unlearning and retention, closely approximating retraining.
@@
Consistent Performance Gains Across Metrics. Our framework consistently outperforms the vanilla baselines across both Forget Quality and Model Utility on both Phi-1.5 and LLaMA2-7B in Table 1. On Phi-1.5, our framework consistently outperforms all vanilla baselines across both Forget Quality and Model Utility. Notably, for instance, under the 5% setting, integrating NPO+GD into our framework significantly elevates the Forget Quality from 0.0878 to 0.7934, while simultaneously improving Model Utility from 0.4158 to 0.4259. Moreover, the ROUGE-L score on the retain set increases from 0.3942 to 0.4147, demonstrating EGUP’s ability to maintain strong generalization while achieving effective unlearning. Similar trends are observed on the LLaMA2-7B model, confirming the generality and robustness of EGUP across architectures and scales. Under the 5% unlearning setting, EGUP(NPO+GD) improves the Forget Quality from 7.54e-5 to 0.9647, while also enhancing Model Utility from 0.5905 to 0.6202 and the Retain ROUGE-L from 0.5681 to 0.7926. Importantly, consistent gains are observed across all other unlearning ratios as well. For example, under the 10% setting, EGUP(NPO+GD) achieves a threefold increase in Forget Quality (0.3417 vs. 5.02e-10) while preserving substantially more useful knowledge, as evidenced by the higher Retain ROUGE-L (0.7266 vs. 0.4815) and Model Utility (0.6315 vs. 0.5608).
@@
Consistent Gains Across Optimizers. Such consistent improvements are observed across all optimizers considered (e.g., GA, GD, NPO, and NPO+GD). Although GA performs less favorably than NPO, and GD underperforms compared to NPO+GD, the integration of all optimizers into our framework leads to noticeable gains. Notably, even though the trade-off between F.Q. and M.U. for GA under our framework remains inferior to the vanilla NPO baseline, it still benefits from our enhancements. We attribute this to the limited potential of the GA optimizer itself. In contrast, the combination of NPO+GD within our framework (denoted as EGUP(NPO+GD)) achieves forget quality that closely approximates that of the retrained model, demonstrating our framework’s strong capability in precise and efficient unlearning.
@@
EGUP framework requires only the average retain embedding yet outperforms methods that use the full retain dataset. The EGUP framework enables even weaker optimizers to outperform stronger baselines by incorporating lightweight, entanglement guidance. A notable example is EGUP(GA), which builds upon GA and requires only the average embedding of the retain dataset. In contrast, GD depends on access to the full original retain dataset. Remarkably, EGUP(GA) consistently outperforms GD across multiple unlearning settings. Under the 5% unlearning setting, EGUP(GA) achieves a higher model utility (0.4261 vs. 0.4047) than GD, while also slightly improving forget quality (1.12e-5 vs. 1.39e-6). This demonstrates that even coarse-grained retain signals, such as a single average embedding vector, can provide strong regularization to guide unlearning without explicit access to the full retain dataset or gradient-level control. The same trend holds across multiple forgetting ratios. For instance, under 10% forgetting, EGUP(GA) boosts utility from 0.3878 to 0.4106, overtaking GD’s 0.4032 again, using only a minimal representation of the retain distribution.
EGUP Consistently Outperforms Baselines Across Phi and LLaMA2-7B. Figure 6 reports unlearning performance on both Phi and LLaMA2-7B across the TOFU benchmarks. Across the two model families, EGUP achieves a better trade-off between forget quality and model utility than alternative methods. The empirical-loss-based heuristic (Huang et al., 2024) is unstable: relative to the vanilla baseline it sometimes yields better unlearning performance but at other times performs worse, indicating high sensitivity to sample-level loss patterns and resulting variance across models and forgetting ratios. SatImp (Yang et al., 2025) tends to preserve model utility well on both Phi and LLaMA, but its forget quality is notably weaker, suggesting incomplete removal of targeted knowledge.
@@
We further illustrate the effectiveness of EGUP by visualizing, in Figure 7 the per-sample absolute loss gap between each unlearned model and the retrained model on both the forget and retain datasets. The top row shows results on the forget dataset, where we observe that models incorporating entanglement guided unlearning (EGU) consistently produce smaller loss gaps compared to their counterparts. EGU better approximate the behavior of retraining in removing memorized knowledge. The bottom row presents results on the retain dataset. Notably, methods built upon our entanglement-guided framework consistently yield loss patterns that more closely align with the retrained model, compared to their non-entanglement counterparts. This indicates that by assigning differentiated unlearning effort based on sample entanglement, our approach enhances unlearning precision while significantly mitigating undesired unlearning of the retain knowledge, thus achieving a more stable and utility-preserving unlearning process.
Figure 9. Performance comparison across four key metrics under three forgetting rates. EGUP(NPO+GD) consistently outperforms the other methods across all metrics.
Figure 9 compares EGUP with baseline approaches on PHI-1.5 and LLaMA2-7B under different forgetting rates (1%, 5%, and 10%). The radar plots evaluate four key metrics: Forget Quality, ForgetRL, Model Utility, and RetainRL. The blue curve denotes the retrain baseline, serving as the upper-bound target, while the orange curve represents the standard NPO+GD method. Introducing entanglement-guided loss reweighting (green) consistently enhances Forget Quality across all settings. Further incorporating the proxy constraint (red) yields notable improvements in Model Utility and brings performance closer to the retrain upper bound across most metrics. Particularly under the 1% forgetting scenario, the red curve almost overlaps with the retrain baseline, indicating that EGUP can closely emulate retraining when the forget ratio is small. As the forget rate increases to 5% and 10%, the performance gap gradually widens, reflecting the increasing difficulty of jointly maintaining forget quality and model utility.
6.2. Ablation Study
To better understand the contribution of each component in our framework, we conduct ablation studies by isolating the two key modules of EGUP. In Section 6.2.1, we evaluate the impact of entanglement-guided unlearning by disabling the proxy constraint. In Section 6.2.2, we assess the effectiveness of the proxy constraint by removing EGU, thereby highlighting its role in regularizing the unlearning process and preventing over-unlearning.
6.2.1. Entanglement-Guided Unlearning Framework
To further evaluate the effect of entanglement-guided loss reweighting in our unlearning framework, we conduct an ablation study on Forget-05, using NPO as the base unlearning method. Figure 8 compares the losses on the forget and retain dataset against the retrained model. It also illustrates how each component, including entanglement-guided loss reweighting, proxy constraint, and their combination, impact the trade-off between unlearning efficacy and model utility.
In Figure 8, each column presents the forget and retain losses under a specific variant of our method, evaluated against the retrained model’s loss as a reference. The first column shows the effect of applying entanglement-guided unlearning (EGU) alone. This mechanism amplifies the influence of highly entangled tokens during unlearning, thereby improving the unlearning performance, evident from the forget loss being closer to that of the retrained model. However, this comes at the cost of increased retain loss, indicating potential over-unlearning or unintended interference with useful knowledge. The second column isolates the impact of proxy regularization based on in-context learning (ICL). Although it contributes little to reducing the forget loss, it substantially improves the retain loss, demonstrating its effectiveness in constraining over-unlearning and preserving model utility. Finally, the third column shows the result of EGUP. This combination achieves the best balance, with both forget and retain losses approaching those of the retrained model. The figure thus highlights the complementary strengths of the two modules and underscores that their integration is critical to achieve precise and utility-preserving unlearning.
6.2.2. ICL-based Proxy Constraint
Figure 4 compares the unlearning dynamics under three forgetting rates (1%, 5%, 10%) between standard gradient difference (dashed) and EGUP_w/o(GD) (solid), across forget loss, retain loss, and penalty loss. We summarize our findings through the following key observations.
Stability in the early unlearning stage. At the initial training steps across all forget rates, both the forget loss and retain loss remain consistently low. This observation indicates that the unlearning process is well-behaved in its early phase. Meanwhile, the model maintains strong performance on the retain dataset, thus avoiding premature over-unlearning.
No penalty activation under mild unlearning. In the 1% forgetting case, no signs of over-unlearning are observed throughout the entire training trajectory. The penalty loss remains zero across all training steps, demonstrating that the test-informed penalty mechanism stays inactive when unnecessary and does not interfere with mild unlearning tasks.
Figure 10. Performance comparison of different unlearning strategies across 1%, 5%, 10% in terms of Forget Quality (left) and Model Utility (right). Black line separates variants without/with proxy constraints. Incorporating test constraint improves utility while maintaining forget quality.